If the third term of a GP is 36 and the 6th term is 243/2 find the first term,the common ratio and the sum of the 8th terms
Answer by richwmiller(17219)

already answered
Tn = t1 * r^(n - 1)
T3=36= t1 * r^(2)
T6 =243/2= t1 * r^(5)
243/2= t1 * r^(5),
36= t * r^(2)
r = 3/2, t1 = 16
S=t1*(1 - r^n)/(1 - r)
S=16*(1 - (3/2)^8)/(1 - 3/2)
S=6305/8
S=788 1/8